How many orbitals in an atom could have these sets of quantum numbers?

n=6 l=3 ml=2
Answer #1

- n= 6 - 3 me = 2 6th shell f orbital one orbital Hence, one orbital of f orbital of 8th shell have n=6, e-3, me=2 so one orb